OpenStudy (anonymous):

a culture started with 4,000 bacteria. After 5 hours, it grew to 4,800 bacteria. Predict how many bacteria will be present after 15 hours

OpenStudy (anonymous):

growth rate=(4800-4000)/5=800/5=160 bacteria/ hour

OpenStudy (anonymous):

after 15 hour bacteria=4000 + growth =4000 + 15*160=4000+ 2400=6400